A place kicker kicks the ball upward with a velocity of 32ft/sec. Ignoring the height of the kicking tee, how long after the ball is kicked does it hit the ground? Use the formula h(t)= -16ft squared + vt where h(t) is the height of the object in feet, v is the ball's initial velocity in ft/sec, and t is time in seconds. What is the highest point reached by the ball?

Answer by Alan3354(69443)

A place kicker kicks the ball upward with a velocity of 32ft/sec. Ignoring the height of the kicking tee, how long after the ball is kicked does it hit the ground? Use the formula h(t)= -16ft squared + vt where h(t) is the height of the object in feet, v is the ball's initial velocity in ft/sec, and t is time in seconds. What is the highest point reached by the ball?
-----------------------
h(t)= -16ft squared + vt
h(t)= -16t^2 + 32t (I think you meant this)
---------
The highest point is the vertex of the parabola, @ t = -b/2a
t = -32/32 = 1 second.
h(1) = -16 + 32 = 16 feet
----------
how long after the ball is kicked does it hit the ground?
It hits the ground when h(t) = 0
h(t)= -16t^2 + 32t = 0
-16t^2 + 32t = 0
16t*(t - 2) = 0
t = 0 (when it's kicked straight up)
t = 2 seconds at impact
============
This guy's out of a job, kicking the ball straight up with a 2 second hang time.